OpenStudy (anonymous):

Find a ⋅ b. (2 points) a = <7, 4>, b = <3, 5> 41 <10, 9> -1 <21, 20>

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

dot product?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

dot product is a number, not a vector, so skip 2 and 4

OpenStudy (misty1212):

\[7\times 3+4\times 5\]

OpenStudy (anonymous):

its 41 :)

OpenStudy (anonymous):

i have another one can u help ill give medal

OpenStudy (anonymous):

Find a ⋅ b. (2 points) a = 9i + 5j, b = 4i + 3j 51 <13, 8> <36, 15> -21 @misty1212

OpenStudy (misty1212):

same as the last one

OpenStudy (misty1212):

\[9\times 4+5\times 3\]

OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. (2 points) u = <2, -4>, v = <3, -8>

OpenStudy (anonymous):

@misty1212